OpenStudy (anonymous):

find the common difference for the arithmetic sequence 6,11,16,21,...

OpenStudy (anonymous):

this one I would just do 11-6

OpenStudy (anonymous):

YES

OpenStudy (anonymous):

common diff = a2-a1 =11-6 =5

OpenStudy (anonymous):

thank you I think it just confuses me because this is for precal and it seems to easy